Yehle Inc. regularly uses material Y51B and currently has in stock 458 liters of the material for which it paid $2,620 several weeks ago. If this were to be sold as is on the open market as surplus material, it would fetch $5.22 per liter. New stocks of the material can be purchased on the open market for $5.92 per liter, but it must be purchased in lots of 1,000 liters. You have been asked to determine the relevant cost of 650 liters of the material to be used in a job for a customer. The relevant cost of the 650 liters of material Y51B is:

Noreen 4e Recheck 2017-16-03

$5,920

$3,848

$3,393

$3,574

question 2)

The Melrose Corporation produces a single product, Product C. Melrose has the capacity to produce 98,000 units of Product C each year. If Melrose produces at capacity, the per unit costs to produce and sell one unit of Product C are as follows:

Direct materials $ 29.80
Direct labor $ 22.60
Variable manufacturing overhead $ 17.20
Fixed manufacturing overhead $ 21.00
Variable selling expense $ 14.80
Fixed selling expense $ 9.40


The regular selling price of one unit of Product C is $132.80. A special order has been received by Melrose from Moore Corporation to purchase 5,000 units of Product C during the upcoming year. If this special order is accepted, the variable selling expense will be reduced by 75%. Total fixed manufacturing overhead and fixed selling expenses would be unaffected except that Melrose will need to purchase a specialized machine to engrave the Moore name on each unit of product C in the special order. The machine will cost $9,300 and will have no use after the special order is filled. Assume that direct labor is a variable cost.

Assume that Melrose expects to sell 88,000 units of Product C to regular customers next year. At what selling price for the 5,000 units would Melrose be economically indifferent between accepting and rejecting the special order from Moore?

rev: 11_01_2016_QC_CS-68231

$95.66

$97.16

$75.16

$73.66
Answer #1

Since the material is regularly used, relevant cost is Purchase cost

= 650*5.92

= $3,848

Excess material will be sued for regular purposes

Since there is spare capacity, relevant costs are

Direct material

       149,000

Direct labor

       113,000

Variable manufacturing overhead

          86,000

Variable selling expenses

          18,500

Cost of machine

            9,300

Relevant cost

       375,800

Indifference price per unit = 375,800/5000 = $75.16